OpenStudy (anonymous):

Part 1: Show the work to solve |5x – 2| 8

OpenStudy (anonymous):

Is the problem supposd to look like 5x -2 = 8?

OpenStudy (jiteshmeghwal9):

well the absolute value is always positive so 5x-2=8

OpenStudy (cwrw238):

there another possibility 5x - 2 = -8

OpenStudy (jiteshmeghwal9):

-(5x-2)=-(8)

OpenStudy (jiteshmeghwal9):

its another possibility

OpenStudy (anonymous):

well its 5x + 2 = 8 if absolute value is always positive

OpenStudy (jiteshmeghwal9):

subtract 2 from both sides & then divide 5 both sides

OpenStudy (cwrw238):

absolute value is always positive bu whats inside the | | can be negative

OpenStudy (jiteshmeghwal9):

becoz let us take |-6|=-(-6)=6

OpenStudy (jiteshmeghwal9):

& if it is a unknown number than we should multiply this to (-)

OpenStudy (cwrw238):

i wonder where hello has gone....

OpenStudy (jiteshmeghwal9):

so this question has two answer

OpenStudy (cwrw238):

yes

OpenStudy (jiteshmeghwal9):

hello! is offline now

OpenStudy (jiteshmeghwal9):

so its answers are 5x-2=8 -(5x-2)=-(8)

OpenStudy (jiteshmeghwal9):

solving these for 'x' will give us answer @hello02713 :)

OpenStudy (cwrw238):

5x - 2 = 8 5x = 10 x = 2

OpenStudy (cwrw238):

your second equation is incorrect

OpenStudy (jiteshmeghwal9):

why ?

OpenStudy (jiteshmeghwal9):

sorry

OpenStudy (jiteshmeghwal9):

-(5x-2)=8

OpenStudy (cwrw238):

it should be either - (5x - 2) = 8 yes